Llulate sodium fluoride after the addition of 0.04 mol of HCl to 1 L. of this solution. Assume no change in volume. K. 7.1*10 the pH of a buffer solution made from 0.30 M hydrofluoric acid and 0.70 5. Calculate the pH of a 0.04 M HCI solution. Compare to the pH found in problem 4 6. Consider the following weak acids and their K values: Acetic acid Phosphoric acid Hypochlorous acid K1.8 x 10-5 Ka 7.5 x 10-3 K 3.5 x10-8 You are to prepare buffers at pH 2.8, 4.5, and 7.5. In the 2ad column write the formulas for the respective weak acid-conjugate base buffer system that is the best choice for each pH from the listed acids? Explain your reasoning in the 3rd column. Rational pH 2.8 4.5 7.5 weak acid-conjugate base combination
Answer #1

4)

Ka = 7.1 x 10^-4

pKa = -log Ka

pKa = -log (7.1 x 10^-4)

pKa = 3.15

pH    = pKa + log [NaF - C / HF + C]

pH = 3.15 + log (0.70 - 0.04 / 0.30 + 0.04)

pH = 3.45

5)

HCl concentration = 0.04 M

[H+] = 0.04 M

pH = -log [H+]

pH = -log (0.04)

pH = 1.40

this pH value is very less compare to problem number 4 pH value . because HCl is strong acid and HF and NaF mixture is weak